# FRIDA to the Rescue! Analyzing Synthetic Data Effectiveness in Object-Based Common Sense Reasoning for Disaster Response

### v2

@@ Abstract
During Human Robot Interactions in disaster relief scenarios, Large Language Models (LLMs) have the potential for substantial physical reasoning to assist in mission objectives. However, these capabilities are often found only in larger models, which are frequently not reasonable to deploy on robotic systems. To meet our problem space requirements, we introduce a dataset and pipeline to create Field Reasoning and Instruction Decoding Agent (FRIDA) models. In our pipeline, domain experts and linguists combine their knowledge to make high-quality few-shot prompts used to generate synthetic data for fine-tuning. We hand-curate datasets for this few-shot prompting and for evaluation to improve LLM reasoning on both general and disaster-specific objects. We concurrently run an ablation study to understand which kinds of synthetic data most affect performance. We fine-tune several small instruction-tuned models and find that ablated FRIDA models only trained on objects’ physical state and function data outperformed both the FRIDA models trained on all synthetic data and the base models in our customized evaluation. We demonstrate that the FRIDA pipeline is capable of instilling physical common sense with minimal data.
@@
Mollie Shichman1, Claire Bonial2, Austin Blodgett2, Taylor Hudson3, Francis Ferraro4, Rachel Rudinger1 1 University of Maryland, College Park, 2 Army Research Lab 3 Oak Ridge Associated Universities, 4 University of Maryland, Baltimore County mshich@umd.edu, claire.n.bonial.civ@army.mil, ferraro@umbc.edu, rudinger@umd.edu
@@ 1 Introduction
Which of the following would be most dangerous if it collapsed? This question, as seen in Figure 1, is fairly trivial for humans to answer, but requires several types of semantic knowledge. First, one must know which objects are capable of collapsing. One must also know the general size of these items and the their other functions to fully assess the danger the item poses. A collapse is also a change of state that fundamentally shifts the use of these objects; a collapsed chair could be more likely to cut or scrape someone, but it could also mean the chair can now be carried if the chair folds. All of this knowledge is needed to answer this question, and all of it is embedded in our semantic understanding of objects that can cause danger and objects that can collapse, both intentionally and unintentionally.
@@
To answer our research question, we first test the effectiveness of fine-tuning smaller models on disaster relief data. However, available data proved to be an additional constraint. Most publicly available data on disasters is social media-based reactions Godinho (2024), which do not pertain much to our subdomain of disaster relief efforts. Furthermore, the specific knowledge (and to a lesser extent, the general knowledge) required for each mission varies by disaster. For example, after an earthquake, a robot needs to find survivors, while after a chemical spill, a robot needs to sample the environment for hazardous materials. Therefore, we need a method for generating training data for specific disasters, and we need to evaluate which data are most effective at improving robot performance.
@@
To investigate which synthetic data most influenced model performance, we ran an ablation study where we fine-tuned the same small LLMs on subsets of our synthetic data corresponding to specific types of object-based reasoning. We call these resulting models the ablated FRIDA (aFRIDA) models. We found that aFRIDA models trained on general semantics and physical common sense had stronger overall performances than models trained on only domain-specific knowledge. Additionally, the best performing aFRIDA models scored better than their corresponding base models and FRIDA models trained on the entire synthetic dataset. We posit that FRIDA succeeds in improving object-related general common sense, but that small LLMs struggle with disaster-specific equipment usage
@@
An anonymous github containing code and a complete walkthrough of the FRIDA pipeline are currently available. [^2] https://anonymous.4open.science/r/FRIDA-0215-D02A/
@@ 2.1 LLMs Reasoning about the World
One resource for improving LLM understanding of an object’s functions, also known as the object’s affordances, is Adak et al. (2024), who curate a dataset of naturally occurring sentences and corresponding images. They then transform them into inference, probing, and masking tasks for LLMs and Visual Language Models (VLMs). Their evaluation shows that VLMs do not have straightforward understandings of object affordances, but few-shot fine-tuning improves LLM and VLM performance on identifying object affordances.
@@ 2.2 Disaster Work and Natural Language Processing
Godinho (2024) completed a systematic search and analysis of over 100 peer-reviewed papers relating to Natural Language Processing (NLP) tools being applied to disasters. 85 of the 107 papers found were for analyzing social media, and the majority of papers focused on sentiment analysis, text classification, and information extraction tasks. Both the data sources and NLP tasks do not have a clear parallel with our objective.
@@ 2.3 Synthetic Data Generation
This paper is inspired in particular by the pipeline used in Wang et al. (2022), who hand crafted 175 “seed” instructions. These seed instructions were used for 8-shot prompting of GPT’s text-davinci-001 model to generate more than 50,000 instructions for a generic and ungrounded AI assistant. These synthetic instructions were then used to fine-tune text-davinci-001. The authors found that their method and resulting fine-tuned model performed comparably to OpenAI’s GPTInstruct Wang et al. (2022). Taori et al. (2023) innovated on Wang et al. (2022) by fine-tuning a separate, smaller language model with a different architecture, as opposed to fine-tuning on the same model that generated the data. They subsequently found that their resulting model’s answers were rated as highly as GPT’s text-davinci-003.
@@ 3.1 FRIDA Seed Data
We developed an expert-in-the-loop pipeline to generate high-quality seed data that leverage expertise on both disaster-relief and semantics. The purpose of this pipeline is to enable quick and efficient fine-tuning of small LLMs to be capable of critical reasoning in specific disaster environments. The details of this pipeline are described in a previous work, here we provide a brief overview. We developed a series of templates that can be filled in with vocabulary from an affordance ontology based on the Rich Event Ontology Kazeminejad et al. (2018). This affordance ontology is extended to serve as an ontology of disaster-related objects and their functionalities, as defined by the objects’ PropBank semantic roles labels Palmer et al. (2005).
@@
Although some related work leverages the same seed sentences used for generating synthetic data to also evaluate the data Wang et al. (2022), we developed a separate evaluation to ensure that our evaluation was not unfairly biased. The seed and evaluation instructions include multiple choice answers, enabling more efficient evaluation and comparison of models.
@@ 3.2 Synthetic Dataset Generation and Analysis
The dataset we use in this work focused on search and rescue operations in the aftermath of the Turkey-Syria Earthquake Arranz et al. (2023). We had 26 templates grouped into 8 categories based on the type of knowledge they query (for all categories and examples, see Table 4). For each template, we made 5 seed instructions for synthetic data generation (130 total instructions) and at least 4 evaluation instructions (119 examples).
@@ 3.3 FRIDA Model construction
We used our synthetic dataset to fine-tune the 1 Billion, 3 Billion, and 8 Billion parameter Instruct models from the LLaMa 3 herd Team (2024b) as well as the Mistral AI’s Ministral 8B Instruction tuned model Team (2024c). We chose to use the LLaMa suite due to it having multiple small instruction tuned models of different sizes, with decently strong performanceTeam (2024b). We chose Ministral 8B to serve as a comparison, since it is trained with sliding window attention, unlike the LLaMa models trained with full attentionTeam (2024c). Additionally, Ministral 8B was released after the LLaMa 3 herd and outperformed the LLaMa models on a variety of metrics Team (2024c). We chose to fine-tune the instruct variations of these models because our task is based in answering questions.
Of the four fine-tuned models, the strongest fine-tuned model performance on our evaluation was from models trained on Ministral 8B. We hypothesize that this is due to the architectural differences between Mistral AI and Meta AI models. Specifically, sliding attention could be helpful in focusing the model’s attention on the instruction content instead of the multiple choice answers. Additionally, Ministral 8B’s sliding attention mechanism is more memory and time efficient, making it more practical for deployment on a robot Team (2024c). As such, we focused our analysis on FRIDA and aFRIDA models based on Ministral 8B, since they are the most conceivable models to work in a robotic system in the near term. Results for the LLaMa models can be found in our anonymous github. Fine-tuning specifics can be found in Appendix section B.
@@ 3.4 Evaluation
Although we leverage multiple choice questions and answers for evaluation, we required a less rigid method than exact match so that formatting errors (e.g. writing “A” instead of “A)”, or forgetting punctuation) would have less impact. Thus, we used SemScore Aynetdinov and Akbik (2024); Geronimo and Lera (2024), which is a scoring metric that uses cosine similarity to compare a model’s embedding vectors of the gold standard and FRIDA responses.
@@ 4 Results
Model SemScore Accuracy (%)
@@
aFRIDA Accuracy (%),
Fine tuning Data Subset all evaluation data
@@ 5 Discussion & Error Analysis
We suspect that the reason “aFRIDA: earthquake” and “aFRIDA: specialized equipment” did not similarly improve performance is that our synthetic data for the more specific objects and tasks tended to be longer and have lower ROUGE scores. These data therefore had more diversity. The sample size of the Earthquakes and Specialized Equipment synthetic data subsets may have been too small for the model to be effectively biased by fine tuning. Conversely, larger models may have ingested operators’ manuals for specialized equipment, facilitating parroting answers for questions on this topic. We note that our research highlights the general difficulty of analyzing the precise effects of fine-tuning given opaque pre-training data.
@@
    “Choose the biggest of a given set of objects in terms of your own common sense”
@@ 5.1 Future Work
There are several ways we can further improve the FRIDA pipeline. We want to improve our prompting for synthetic data to make them less trivial. We can refine and expand our less technical templates. By adding different phrasing, we hope to make our synthetic data more reflective of real world natural language. We also hope implementing the strategies in other work Ge et al. (2024); Ding et al. (2023); Mukherjee et al. (2023) for diversifying synthetic data will improve generation quality and efficiency. Finally, we plan to test the pipeline on other domains with experts to help us refine our process.
@@ Appendix A Categories and Descriptions
Choose the truck from the list. coupe, minivan, 18 wheeler, sedan, ATV
@@
Primary and Secondary Object Facts Where Object Found, Objects in Location, Secondary Uses Hey, which of the following can be used as a lever? art, motorcycle, picture, dvd, broom 15
@@ Appendix D Licenses
We used TRL von Werra et al. (2020) under the Apache License. SemScore Geronimo and Lera (2024) implements the MIT license, and the LLaMa models were used after author agreement to the LLaMa 3.1 and 3.2 Community License Agreement Team (2024b). Ministral 8B Instruct was used under the Mistral Research License Jiang et al. (2023). We will release our code, datasets, and FRIDA models under a open-source license which will be chosen upon conference acceptance.

### v3

@@ Abstract
During Human Robot Interactions in disaster relief scenarios, Large Language Models (LLMs) have the potential for substantial physical reasoning to assist in mission objectives. However, these reasoning capabilities are often found only in larger models, which are not currently reasonable to deploy on robotic systems due to size constraints. To meet our problem space requirements, we introduce a dataset and pipeline to create Field Reasoning and Instruction Decoding Agent (FRIDA) models. In our pipeline, domain experts and linguists combine their knowledge to make high-quality, few-shot prompts used to generate synthetic data for fine-tuning. We hand-curate datasets for this few-shot prompting and for evaluation to improve LLM reasoning on both general and disaster-specific objects. We concurrently run an ablation study to understand which kinds of synthetic data most affect performance. We fine-tune several small instruction-tuned models and find that ablated FRIDA models only trained on objects’ physical state and function data outperformed both the FRIDA models trained on all synthetic data and the base models in our evaluation. We demonstrate that the FRIDA pipeline is capable of instilling physical common sense with minimal data.
@@
Mollie Shichman1, Claire Bonial2, Austin Blodgett2, Taylor Pellegrin3, Francis Ferraro4, Rachel Rudinger1 1University of Maryland, College Park, 2Army Research Lab 3Oak Ridge Associated Universities, 4University of Maryland, Baltimore County mshich@umd.edu, claire.n.bonial.civ@army.mil, ferraro@umbc.edu, rudinger@umd.edu
@@ 1 Introduction
Which of the following would be most dangerous if it collapsed? This question, as seen in Figure 1, is fairly trivial for humans to answer, but requires several types of semantic knowledge. One must know the general size of these items and their other functions to fully assess the danger the item poses. A collapse is also a change of state that fundamentally shifts the use of these objects; a collapsed chair could be more likely to cut or scrape someone, but it could also mean the chair can now be carried if the chair folds. All of this knowledge is needed to answer this question, and all of it is embedded in our semantic understanding of objects that can cause danger and objects that can collapse, both intentionally and unintentionally.
@@
To answer our research question, we first tested the effectiveness of fine-tuning smaller models on disaster relief data. However, available data proved to be an additional constraint. Most publicly available data on disasters is social media-based reactions Godinho (2024), which do not pertain much to our subdomain of disaster relief efforts. Furthermore, the specific knowledge (and to a lesser extent, the general knowledge) required for each mission varies by disaster. For example, after an earthquake, a robot needs to find survivors, while after a chemical spill, a robot needs to sample the environment for hazardous materials. Therefore, we need a method for generating training data for specific disasters, and we need to evaluate which data are most effective at improving robot performance.
@@
To investigate which synthetic data most influenced model performance, we ran an ablation study where we fine-tuned the same small LLMs on subsets of our synthetic data corresponding to specific types of object-based reasoning. We call these resulting models the ablated FRIDA (aFRIDA) models. We found that aFRIDA models trained on general semantics and physical common sense had stronger overall performances than models trained on only domain-specific knowledge. Additionally, the best performing aFRIDA models scored better than their corresponding base models and FRIDA models trained on the entire synthetic dataset. We posit that FRIDA succeeds in improving object-related general common sense, but that small LLMs struggle with disaster-specific equipment usage.
@@
Our datasets, code, and a complete walkthrough of the FRIDA pipeline are currently available. [^2] https://github.com/mshich1/FRIDA/
@@ 2.1 LLMs Reasoning about the World
One resource for improving LLM understanding of an object’s functions, also known as the object’s affordances, is Adak et al. (2024), who curate a dataset of naturally occurring sentences and corresponding images. They then transform them into inference, probing, and masking tasks for LLMs and Visual Language Models (VLMs). Their evaluation shows that VLMs do not have straightforward understandings of object affordances, but few-shot fine-tuning improves LLM and VLM performance on identifying object affordances. This work focuses on building a stronger basis in LLMs to improve these downstream tasks, as well as understand which data are most important for a robot’s success.
@@ 2.2 Disaster Work and Natural Language Processing
Godinho (2024) completed a systematic search and analysis of over 100 peer-reviewed papers relating to Natural Language Processing (NLP) tools being applied to disasters. 85 of the 107 papers found were analyzing social media, and the majority of papers focused on sentiment analysis, text classification, and information extraction tasks. Both the data sources and NLP tasks do not have a clear parallel with our objective.
@@ 2.3 Synthetic Data Generation
This paper is inspired in particular by the pipeline developed by Wang et al. (2022), who hand crafted 175 “seed” instructions. These seed instructions were used for 8-shot prompting of GPT’s text-davinci-001 model to generate more than 50,000 instructions for a generic and ungrounded AI assistant. These synthetic instructions were then used to fine-tune text-davinci-001. The authors found that their method and resulting fine-tuned model performed comparably to OpenAI’s GPTInstruct Wang et al. (2022). Taori et al. (2023) innovated on Wang et al. (2022) by fine-tuning a separate, smaller language model with a different architecture, as opposed to fine-tuning on the same model that generated the data. They subsequently found that their resulting model’s answers were rated as highly as GPT’s text-davinci-003.
@@ 3.1 FRIDA Seed Data
We developed an expert-in-the-loop pipeline to generate high-quality seed data that leverage expertise on both disaster-relief and semantics. The purpose of this pipeline is to enable quick and efficient fine-tuning of small LLMs to be capable of critical reasoning in specific disaster environments. The details of this pipeline are described in Shichman et al. (2024), here we provide a brief overview. We developed a series of templates that can be filled in with vocabulary from an affordance ontology based on the Rich Event Ontology Kazeminejad et al. (2018). This affordance ontology is extended to serve as an ontology of disaster-related objects and their functionalities, as defined by the objects’ PropBank semantic roles labels Palmer et al. (2005).
@@
Although some related work leverages the same seed sentences used for generating synthetic data to also evaluate the data Wang et al. (2022), we used this same pipeline to develop a separate and unique evaluation to ensure that our evaluation was not present in any training data. The seed and evaluation instructions include multiple choice answers, enabling more efficient evaluation and comparison of models.
@@ 3.2 Synthetic Dataset Generation and Analysis
The dataset we use in this work focused on search and rescue operations in the aftermath of the Turkey-Syria Earthquake Arranz et al. (2023). We had 26 templates grouped into 8 categories based on the type of knowledge they query as defined by the Generative Lexicon Qualia Pustejovsky and Jezek (2016). For all categories and examples, see Table 4 of Appendix A. For each template, expert annotators hand-made 5 seed instructions for synthetic data generation (130 total instructions) and a minimum of 4 evaluation instructions (119 examples). All resulting instructions were examined by a second author for correctness.
@@ 3.3 FRIDA Model construction
We used our synthetic dataset to fine-tune the 1 Billion, 3 Billion, and 8 Billion parameter Instruct models from the LLaMa 3 herd Team (2024b) as well as the Mistral AI’s Ministral 8B Instruction tuned model Team (2024c). We chose to use the LLaMa suite due to it having multiple small instruction tuned models of different sizes, with strong performance Team (2024b). We chose Ministral 8B to serve as a comparison, since it is trained with sliding window attention, unlike the LLaMa models trained with full attention Team (2024c). Additionally, Ministral 8B was released after the LLaMa 3 herd and outperformed the LLaMa models on a variety of metrics Team (2024c). We chose to fine-tune the instruct variations of these models because our task is based on answering questions. All models were trained with the performance enhanced fine-tuning model LoRA Hu et al. (2021) with full precision.
Of the four fine-tuned models, the strongest fine-tuned model performance on our evaluation was from models trained on Ministral 8B. We hypothesize that this is due to the architectural differences between Mistral AI and Meta AI models. Specifically, sliding attention could be helpful in focusing the model’s attention on the instruction content instead of the multiple choice answers. Additionally, Ministral 8B’s sliding attention mechanism is more memory and time efficient, making it more practical for deployment on a robot Team (2024c). As such, we focused our analysis on FRIDA and aFRIDA models based on Ministral 8B, since they are the most conceivable models to work in a robotic system in the near term. Results for the LLaMa models can be found in our github. Fine-tuning specifics can be found in Appendix B.
@@ 3.4 Evaluation
As described in section 3.2, we used the same pipeline for creating seed data to create a custom evaluation, with at least four evaluation questions per template for a total of 119 evaluation instructions.
Although we leverage multiple choice questions and answers for evaluation, we required a less rigid method than exact match so that formatting errors (e.g., writing “A” instead of “A)”, or forgetting punctuation) would have less impact. Thus, we used SemScore Aynetdinov and Akbik (2024); Geronimo and Lera (2024), which is a scoring metric that uses cosine similarity to compare a model’s embedding vectors of the gold standard and FRIDA responses.
@@ 4 Results
Model SemScore
Accuracy (%)
@@
Ablated Model SemScore
Fine-Tuning Data Subset Accuracy (%)
@@ 5 Discussion & Error Analysis
We suspect that the reason “aFRIDA: earthquake” and “aFRIDA: specialized equipment” did not similarly improve performance is that our synthetic data for the more specific objects and tasks tended to be longer and have lower ROUGE scores. These data therefore had more diversity. The sample size of the Earthquakes and Specialized Equipment synthetic data subsets may have been too small for the model to be correctly biased by fine tuning. Conversely, larger models may have ingested operators’ manuals for specialized equipment, facilitating parroting answers for questions on this topic. We note that our research highlights the general difficulty of analyzing the precise effects of fine-tuning given opaque pre-training data.
@@
    “Choose the biggest of a given set of objects in terms of your own common sense.”
@@
When it came to reasoning about the complex equipment used, error analysis revealed that both vanilla and fine-tuned models scored perfectly when asked to choose the correct role for an object in an event. For example:
  1. 4.
    “What role does a helicopter play in the search and rescue process?”
    A) Provide a vantage point to identify heavily damaged areas
    B) Move large vehicles to disaster area
    C) Blow away debris
    D) Warn victims about aftershocks
    E) Blow debris out of the way
    Gold: A) Provide a vantage point to identify heavily damaged areas
    Ministral: A) Provide a vantage point to identify heavily damaged areas
    FRIDA: A) Provide a vantage point to identify heavily damaged areas
The task of choosing the correct object to use for a task proved more challenging. Fine-tuning on related data seemed to unnecessarily bias the model toward choosing the most complicated object, while fine-tuning on unrelated data maintained results. For example:
  1. 5.
    “Select the equipment needed for breaking rubble into smaller pieces after an earthquake.”
    A) axe
    B) pickaxe
    C) hydraulic lift
    D) hard hat
    E) hammer
    Gold: B) pickaxe
    Ministral: B) pickaxe
    FRIDA: C) hydraulic lift
    aFRIDA relative sizes: B) pickaxe
In the most complex reasoning task of ordering steps to complete to use an object, fine-tuning had no clear effect, with all models providing random answers.
  1. 6.
    “The following are two different steps for using a dump truck. Which needs to happen first?
    A) Wait for others to fill the truck bed
    B) open the tailgate
    Gold: B) open the tailgate
    Ministral: B) open the tailgate
    FRIDA: B) open the tailgate
    aFRIDA relative sizes: A) Wait for others to fill the truck bed
    aFRIDA required equipment: A) Wait for others to fill the truck bed
We thus conclude that fine-tuning for required equipment did not effectively bias the models to understand the use cases of these complex objects. At its worst, it incorrectly biases the model to choose complex objects when simpler ones would be more effective.
@@ 5.1 Future Work
There are several ways we can further improve the FRIDA pipeline. We want to improve our prompting for synthetic data to make them less trivial to answer. We can refine and expand our less technical templates. By adding different phrasing, we hope to make our synthetic data more reflective of real world natural language. We also hope implementing the strategies in other work Ge et al. (2024); Ding et al. (2023); Mukherjee et al. (2023) for diversifying synthetic data will improve generation quality and efficiency. We want to explore the impact of using quantized models over full precision models to determine if we can save additional storage space while maintaining reasoning ability. Finally, we plan to test the pipeline on other domains with experts to help us refine our process.
@@ Appendix A Categories and Descriptions
Choose the truck from the list: coupe, minivan, 18 wheeler, sedan, ATV
@@
Primary and Secondary Object Facts Where Object Found, Objects in Location, Secondary Uses Which of the following can be used as a lever? art, motorcycle, picture, dvd, broom 15
@@ Appendix D Licenses
We used TRL von Werra et al. (2020) under the Apache License. SemScore Geronimo and Lera (2024) implements the MIT license, and the LLaMa models were used after author agreement to the LLaMa 3.1 and 3.2 Community License Agreement Team (2024b). Ministral 8B Instruct was used under the Mistral Research License Jiang et al. (2023).